How many atoms are in 5.60 moles of Rb


1
Expert's answer
2021-05-10T01:25:42-0400

N-number of atoms

Na-constant number of Avagadros

V-amount of substance


N=Na*V

N=(6.02*10^23)*5.6mol=33,712*10^23